# Privacy Perspectives and Practices of Chinese Smart Home Product Teams

### v1

@@ 4.1.2 Personal perspectives and practices
Strong commitment to privacy and ethics. Most participants demonstrated a strong commitment to user privacy and ethics, with empathy that guided their decisions and helped them balance ethical considerations, technical implementation, and user privacy protection. For example, P1-EN noted: “If the company plans to develop a product that involves risks of infringing user privacy, I will decline to join the development, even if it results in a poor KPI.” Furthermore, some shared that their own discomfort with being monitored by smart home devices directly influenced their approach to privacy-focused development. These personal experiences shaped their efforts to prioritize user privacy in the products they helped create. For example, P12-PM noted that: “informed user consent, transparent data practices, and voluntary data sharing are key components.”
@@
User ownership and control of data in multi-user smart homes. Participants emphasized the importance of user ownership rights, particularly regarding access to, management of, and control over their data. By granting or revoking permissions, users could determine who had access to their data and when, helping the primary user maintain control over their smart home while balancing convenience with privacy and security. For example, some participants developing smart home cameras, doorbells, and door locks described how their design assigned a “super administrator” role to a single account. P11-PM noted: “The owner can grant permissions to friends, domestic workers, or even strangers. This includes allowing them to activate certain features, like viewing real-time video streaming or accessing stored video clips, but within predefined periods.”
@@
Limited exposure to legal and privacy training. Most participants reported some engagement with legal data privacy and protection training programs, although they noted that the training was often superficial. For instance, P14-EN from a small PE mentioned: “Our internal training focused more on real-life development cases rather than providing in-depth legal knowledge.” Similarly, P6-UX expressed skepticism about the focus of these programs: “These training programs seem to be aimed at protecting the company, not the users.” P13-UX also described the training as generic and not specifically tailored to address the privacy needs of smart home camera users: “We have an online exam on confidentiality or permissions every month, or every one to two months.” Further, many participants relied heavily on their companies’ legal or regulatory team members to handle compliance issues, often viewing legal compliance as separate from the main design, development, and testing phases (see §4.1.1). Interestingly, P18-LE highlighted their team’s proactive role in delivering training sessions: “We often provide training for data compliance and privacy, security laws and regulations […]. Some teams may invite us for smaller and more focused training sessions. All training sessions are recorded for future reference.”
@@
Recommendations to improve privacy practices. Many participants suggested that companies should increase transparency in data practices, ensuring users were clearly informed about data collection, sharing, and deletion. For example, P1-EN, P4-PM, & P17-EN advocated for an “overdesign” strategy: “When designing privacy features involves adapting to changing laws and user needs by going beyond standard compliance, adding stricter rules, better privacy features and disclosures, and improved data handling practices.” While P2-PM and P14-EN mentioned the need for a “top-down” approach for privacy requirements—such as establishing a clear development roadmap—P14-EN stated: “It is essential to involve privacy regulation experts during the initial design and development phases, rather than only during or after testing.” Additionally, some participants highlighted future development challenges related to integrating advanced AI systems into smart home devices. For example, P14-EN, P17-EN, & P24-EN noted that embedding generative AI (e.g., Deepseek) in smart home devices could introduce new privacy concerns, as also reported in the academic literature [142].
@@
Perceptions of Chinese users’ privacy awareness and cultural influences. While several participants recognized the importance of addressing privacy needs in China’s smart home market, most believed that privacy awareness among Chinese users varied significantly, influenced by cultural and regional factors. Some felt that only a small fraction of the population was highly privacy-conscious, actively choosing products with strong privacy features, such as creating guest accounts or using surveillance devices only when away from home. Others thought that user privacy concerns were largely influenced by where they resided, with P25-MS stating: “Residents in more developed areas, such as Shanghai and other first-tier cities, tend to have greater privacy awareness compared to those in less developed regions.” Several participants believed that most Chinese users were relatively unconcerned about privacy risks compared to users in Western countries: “People with higher privacy awareness constitute even less than 1%.” (P14-EN). P11-PM & P14-EN linked this to China’s collectivist vs. Western more individualist culture [56]. P14-EN explained, “The education we receive from a young age emphasizes collectivism, and placing too much emphasis on personal interests is viewed as selfish in our cultural values.” This cultural context led some Chinese product teams to downplay privacy in their product designs, assuming that Chinese users prioritized collective benefits over individual privacy. Further, P19-LE discussed the influence of national laws on individual privacy awareness, stating: “Data compliance regulations in China differ from those in Europe. In China, national security takes precedence in data compliance, leading individuals to prioritize national security over their own data privacy and security.”
@@
Many participants emphasized the necessity of collecting personally identifiable information (PII), such as phone numbers, for account registration to comply with China’s real-name system under the CSL. Since phone numbers are tied to real-world identities, participants believed this helped reduce security risks by allowing law enforcement to trace malicious activity. Additionally, several noted that the widespread use of CCTV and facial recognition in public spaces across China had fostered general acceptance of such technologies, resulting in fewer concerns about their use in domestic settings. This was contrasted with certain Western cultures, where stronger privacy concerns may hinder the adoption of home surveillance devices [127, 77].
@@
Summary. Participants’ personal experiences with smart homes and the discomfort of being monitored had fostered empathy toward users and a desire to balance privacy with functionality. However, many believed that Chinese users prioritized collective benefits over individual privacy, leading them to de-prioritize privacy in product design.
@@ 4.2.1 Development challenges
We outline the challenges participants and their teams faced during development, including limited attention to at-risk groups, difficulty balancing privacy with user experience, and broader constraints such as limited resources and development pressures.
Challenges in designing for at-risk groups. Participants faced significant challenges in gathering requirements from at-risk groups such as children, older adults, and domestic workers. Limited resources prevented participants from conducting direct research with these users, making it difficult to understand and address their specific needs. For example, P17-EN noted that their research prioritized “the majority [of] population in China”, often excluding groups such as survivors of intimate partner violence (IPV). P5-UX and P7-PM pointed to the challenges of recruiting IPV survivors and gaining their trust, as many were hesitant to disclose personal experiences. This made it especially difficult to design effective and inclusive solutions for vulnerable populations. Furthermore, P6-UX described the challenges their team faced when designing smart TVs for elder and disabled users, citing a lack of expertise in understanding the specific privacy concerns of these groups. P4-PM’s team implemented protective measures for elder users: “We use millimeter-wave radars in our cameras that capture only coarse-grained data, such as blurred body shapes, to protect privacy, especially in elderly care homes.”
Challenges in balancing privacy and UX. Although most participants acknowledged the importance of privacy, they also recognized the necessity of collecting data to enhance products, services, convenience, and user experience. This created a development paradox: emphasizing privacy required collecting less user data, which in turn could restrict improvements to the UX. As P24-EN said: “Collecting extensive user data is essential for improving our products; without it, providing personalized and optimized UX becomes more difficult.” Some participants mentioned that many Chinese users were willing to trade privacy for convenience, posing a significant challenge for product teams aiming to balance the two. For example, P9-UX noted that long privacy policies could undermine usability, while P11-PM protected user privacy through robust security measures: “For security reasons, we implemented complex password requirements to access cameras. However, we received considerable negative feedback in the app store, and many users found the process cumbersome and not user-friendly.”
@@ 4.2.2 Power Dynamics
Companies vs. companies. Some participants highlighted the complex and often challenging relationships between small and large tech companies, particularly when integrating into tech giants’ ecosystems, which imposed strict compliance and quality standards. P10-PM and P12-PM elaborated on these challenges: “Joining a tech giant’s smart home ecosystem forces small firms to meet strict standards that may conflict with user preferences and hinder integration. Without specialized expertise, these firms often face compatibility issues and must invest extra effort to align with larger platforms.” However, P18-LE argued that integration into tech giant’s ecosystems often compelled smaller firms to comply with strict standards not only due to regulatory requirements but also due to concerns about maintaining their reputation: “Compliance from these small companies is essential not only for meeting legal standards, but also for maintaining the reputation of larger players.”
@@
Companies vs. product teams. Many participants emphasized that internal power dynamics within their companies often restricted the autonomy of product teams, making it difficult to prioritize or implement privacy measures. Senior management, marketing, or business units frequently had greater influence, leading to decisions that favored marketability or speed over user privacy. For instance, P15-EN noted: “This decision-making process often compelled us to prioritize company objectives over user-centered design and privacy-focused innovations.” Further, some participants indicated that internal policies and corporate culture played a major role in how privacy measures were implemented.
@@
Users vs. users. Many participants pointed out the complex power dynamics present in multi-user smart homes, especially tensions between parental monitoring and children’s privacy, hosts and guests, as well as employers and domestic workers. For example, P3-PM, P12-PM, and P26-MS described conflicts emerging from the use of smart home cameras in parent-teen relationships, conflicts shaped by Confucian values that stress filial piety and deference to parental authority, as P3-PM explained: “Many Chinese parents buy these cameras primarily to help with homework and monitor their children’s posture while studying […]. Children often find it difficult to challenge the decisions of their parents, […] In Chinese families, questioning the choice of a parent is often seen as disrespectful.” In addition, in Chinese culture, where hospitality is highly valued, guests may not be aware of being monitored or feel uneasy knowing that smart devices are present, which can be perceived as a breach of trust and courtesy. Additionally, P9-UX, P18-LE, and P20-LE noted the growing use of smart home devices to monitor domestic workers—often migrants from rural areas, as in [51]—for security or deterrence purposes, raising ethical concerns about their privacy: “Privacy and security settings are typically designed to protect device owners, often overlooking the comfort of domestic workers who lack the power to consent to being monitored by their employers’ surveillance cameras.” Further, P12-PM observed that purchasing and installing smart home devices often mirrored existing household gender roles, similar to findings reported in [9]: “Gender dynamics often shape purchasing decisions in typical Chinese households: when a woman suggests buying a device, she may seek input from other family members, but if a man—usually the father or husband—disagrees, the purchase often does not happen.”
@@
Law enforcement vs. companies. Some participants noted that their companies usually required legal documents, such as a search warrant, to authorize police access to user data, in accordance with PIPL (Articles 34, 35). P5-UX added that they generally avoided granting access to data unless it was essential for an investigation, sometimes obtaining the consent of the homeowner beforehand. However, some participants noted that certain companies allowed police access to all user registration and facial data collected by smart home cameras in specific emergency situations. For example, P13-UX highlighted: “Government or public security departments have special privileges to access user facial data, ID, names, and other personal information. They have the authority to force us to provide this data and can also retrieve it through internal channels, bypassing the restrictions our teams put in place.”
@@ 4.3.1 Technical interventions
To mitigate privacy concerns and comply with regulations, many participants described anonymizing or blurring image data prior to processing. As P9-UX explained, “We don’t capture real images directly—sensitive data, like nude bodies, is processed by blurring privacy zones detected by smart cameras.” Furthermore, several company guidelines imposed strict procedures for managing data, especially when it involved sensitive items such as biometric records. For example, P11-PM and P24-EN noted that camera footage was encrypted prior to transmission, and their team built specialized decoders to retrieve those recordings. P9-UX and P17-EN described employing a data-minimization approach, with P9-UX explaining: “We only send and save posture data, like simplified 2D images resembling stick figures, instead of actual scene details.” Moreover, many participants mentioned that their companies followed data deletion practices, often in accordance with regulatory mandates. For example, P4-PM, P16-PM, and P19-LE noted that Chinese regulatory bodies required companies to respect users’ rights to have their data erased. P5-UX further noted: “Given the large volume of data and the high cost of storage, we typically delete video footage after it has been streamed in real time, unless users have subscribed to a storage service.” P11-PM, P21-EN, and P22-EN also mentioned that users had the option to delete their cloud-stored data within 30 days, after which the system would automatically remove it.
Most participants highlighted that their companies imposed strict controls on sensitive user information, allowing only authorized personnel to access it. For example, P12-PM explained: “Our software developers have different levels of access. For example, a developer with a user account can view linked devices, whereas database and server developers are limited to operational queries, such as tracking user numbers or feature usage, without accessing individual user data. Once a feature is deployed, it is controlled by the user, and other developers no longer need to maintain or access it.” However, some participants, especially UX designers, expressed a lack of clarity regarding how smart home devices managed and processed data, often citing limited access and strict company policies as the reason. As P6-UX stated: “We often lack a full understanding of data collection and processing because our design team operates separately from the compliance team.” P18-LE and P23-LE also underscored this issue from the compliance team’s perspective, noting that limited cross-team communication.

### v2

@@ 4.1.2 Personal perspectives and practices
Strong commitment to privacy and ethics. Most participants demonstrated a strong commitment to user privacy and ethics, with empathy that guided their decisions and helped them balance ethical considerations, technical implementation, and user privacy protection. For example, P1-EN noted: “If the company plans to develop a product that poses risks of infringing on user privacy, I will decline to participate in its development, even if it negatively affects my KPI.” Furthermore, some shared that their own discomfort with being monitored by smart home devices directly influenced their approach to privacy-focused development. These personal experiences shaped their efforts to prioritize user privacy in the products they helped create. For example, P12-PM noted that: “informed user consent, transparent data practices, and voluntary data sharing are key components.”
@@
User ownership and control of data in multi-user smart homes. Participants emphasized the importance of user ownership rights, particularly regarding access to, management of, and control over their data. By granting or revoking permissions, users could determine who had access to their data and when, helping the primary user maintain control over their smart home while balancing convenience with privacy and security. For example, some participants developing smart home cameras, doorbells, and door locks described how their design assigned a “super administrator” role to a single account. P11-PM noted: “The owner can grant permissions to friends, domestic workers, or even strangers. These permissions may include activating certain features, such as viewing real-time video streams or accessing stored video clips, but only within predefined time periods.”
@@
Limited exposure to legal and privacy training. Most participants reported some engagement with legal data privacy and protection training programs, although they noted that the training was often superficial. For instance, P14-EN from a small PE mentioned: “Our internal training focused more on real-life development cases rather than providing in-depth legal knowledge.” Similarly, P6-UX expressed skepticism about the focus of these programs: “These training programs seem to be aimed at protecting the company, not the users.” P13-UX also described the training as generic and not specifically tailored to address the privacy needs of smart home camera users: “We have an online exam on confidentiality or permissions every month, or every one to two months.” Further, many participants relied heavily on their companies’ legal or regulatory team members to handle compliance issues, often viewing legal compliance as separate from the main design, development, and testing phases (§4.1.1). Interestingly, P18-LE highlighted their team’s proactive role in delivering training sessions: “We often provide training for data compliance and privacy, security laws and regulations […]. Some teams may invite us for smaller and more focused training sessions. All training sessions are recorded for future reference.”
@@
Recommendations to improve privacy practices. Many participants suggested that companies should increase transparency in data practices, ensuring users were clearly informed about data collection, sharing, and deletion. For example, P1-EN, P4-PM, and P17-EN advocated for an “over-design” strategy: “Designing privacy features involves adapting to evolving laws and user needs by going beyond standard compliance—implementing stricter rules, enhanced privacy controls and disclosures, and improved data handling practices.” While P2-PM and P14-EN mentioned the need for a “top-down” approach to defining privacy requirements—such as establishing a clear development road-map—P14-EN stated: “It is essential to involve privacy regulation experts during the initial design and development phases, rather than only during or after testing.” Additionally, some participants highlighted future development challenges related to integrating AI into smart home devices. For example, P14-EN, P17-EN, and P24-EN noted that embedding generative AI (e.g., Deepseek) in smart home devices could introduce new privacy concerns, as also reported in the academic literature [127].
@@
Perceptions of Chinese users’ privacy awareness and cultural influences. While several participants recognized the importance of addressing privacy needs in China’s smart home market, most believed that privacy awareness among Chinese users varied significantly, influenced by cultural and regional factors. Some felt that only a small fraction of the population was highly privacy-conscious, actively choosing products with strong privacy features, such as creating guest accounts or using surveillance devices only when away from home. Others thought that user privacy concerns were largely influenced by where they resided, with P25-MS stating: “Residents in more developed areas, such as Shanghai and other first-tier cities, tend to have greater privacy awareness compared to those in less developed regions.” Several participants believed that most Chinese users were relatively unaware of or unconcerned about privacy risks compared to users in Western countries: “People with higher privacy awareness constitute even less than 1%” (P14-EN). Participants (e.g., P11-PM & P14-EN) framed collectivism as one influence among several—such as professional ethics, personal values, and market demands—rather than as the sole driver of their privacy practices. P14-EN explained, “The education we receive from a young age emphasizes collectivism, and placing too much emphasis on personal interests is viewed as selfish in our cultural values.” This cultural context led some Chinese product teams to downplay privacy in their product designs, assuming that Chinese users prioritized collective benefits over individual privacy. Furthermore, P19-LE discussed the influence of national laws on individual privacy awareness, stating: “Data compliance regulations in China differ from those in Europe. In China, national security takes precedence in data compliance, leading individuals to prioritize national security over their own data privacy and security.”
@@
Many participants emphasized the necessity of collecting personally identifiable information (PII), such as phone numbers, for account registration to comply with China’s real-name system under CSL. Since phone numbers are tied to real-world identities, participants believed that this helped reduce security risks by allowing law enforcement to trace malicious activity. Additionally, several noted that the widespread use of CCTV and facial recognition in public spaces across China had fostered general acceptance of such technologies, resulting in fewer concerns about their use in domestic settings. This contrasted with participants’ perceptions of Western contexts, where they believed that heightened privacy concerns could impede the adoption of home surveillance devices [113, 68].
@@
Summary. Participants’ personal experiences with smart homes and the discomfort of being monitored had nurtured empathy for users and the desire to balance privacy with functionality. However, many believed that Chinese users prioritized collective benefits over individual privacy, leading them to de-prioritize privacy in product design. Yet, privacy was negotiated between perceived user attitudes, legal requirements, and socio-cultural norms, with some participants prioritizing it on ethical grounds even when they believed users would not.
@@ 4.2.1 Development challenges
We outline the challenges participants and their teams faced, including limited attention to at-risk groups, difficulty balancing privacy with UX, and broader constraints such as limited resources and development pressures. Some participants also distinguished between data essential to core device functions (e.g., camera footage for incident detection or remote monitoring, access-control logs) and non-essential data (e.g., personalization signals and analytics). For essential data, they emphasized necessity, proportionality, and minimal retention, whereas for non-essential data, they favored opt-in defaults, granular controls, and data minimization.
Challenges in designing for at-risk groups. Participants faced significant challenges in gathering requirements from at-risk groups such as children, older adults, and domestic workers. Limited resources prevented participants from conducting direct research with these users, making it difficult to understand and address their specific needs. For example, P17-EN noted that their research prioritized “the majority population in China,” often excluding groups such as survivors of intimate partner violence (IPV). P5-UX and P7-PM pointed out the challenges of recruiting IPV survivors and gaining their trust, as many were reluctant to disclose personal experiences. This made it especially difficult to design effective and inclusive solutions for vulnerable populations. Furthermore, P6-UX described the challenges their team faced when designing smart TVs for older and disabled users, citing a lack of experience in understanding the specific privacy concerns of these groups. P4-PM’s team implemented protective measures for older adults: “We use millimeter-wave radars in our cameras that capture only coarse-grained data, such as blurred body shapes, to protect privacy, especially in elderly care homes.”
Challenges in balancing privacy and UX. Although most participants acknowledged the importance of privacy, they also recognized the necessity of collecting data to enhance products, services, convenience, and user experience. This created a development paradox: emphasizing privacy required collecting less user data, which, in turn, could restrict improvements to UX. As P24-EN said: “Collecting extensive user data is essential for improving our products; without it, providing personalized and optimized UX becomes more difficult.” Some participants mentioned that many Chinese users were willing to trade privacy for convenience, posing a significant challenge for product teams aiming to balance the two. For example, P9-UX noted that long privacy policies could undermine usability, while P11-PM protected user privacy/security through robust measures: “For security reasons, we implemented complex password requirements to access cameras. However, we received considerable negative feedback in the app store, and many users found the process cumbersome and not user-friendly.”
@@ 4.2.2 Power Dynamics
Companies vs. companies. Some participants highlighted the complex and often challenging relationships between small and large tech companies, particularly when integrating into tech giants’ ecosystems, which imposed strict compliance and quality standards. P10-PM and P12-PM elaborated on these challenges: “Joining a tech giant’s smart home ecosystem forces small firms to meet strict standards that may conflict with user preferences and hinder integration. Without specialized expertise, these firms often face compatibility issues and must invest extra effort to align with larger platforms.” P18-LE argued that integration into tech giants’ ecosystems often compelled smaller firms to comply with strict standards, not only because of regulatory requirements but also out of concern for maintaining their reputation: “Compliance from these small companies is essential not only for meeting legal standards, but also for maintaining the reputation of larger players.”
@@
Companies vs. product teams. Many participants emphasized that the dynamics of internal power within their companies often restricted the autonomy of product teams, making it difficult to prioritize or implement privacy measures. Senior management, marketing, or business units frequently had greater influence, leading to decisions that favored marketability or speed over user privacy. For instance, P15-EN noted: “This decision-making process often compelled us to prioritize company objectives over user-centered design and privacy-focused innovations.” Further, some participants indicated that internal policies and corporate culture played a major role in how privacy measures were implemented.
@@
Users vs. users. Many participants pointed out the complex power dynamics present in multi-user smart homes, especially tensions between parental monitoring and children’s privacy, hosts and guests, as well as employers and domestic workers. For example, P3-PM, P12-PM, and P26-MS described conflicts emerging from the use of smart home cameras in parent-teen relationships, conflicts shaped by Confucian values stressing filial piety and deference to parental authority, as P3-PM explained: “Many Chinese parents buy these cameras primarily to help with homework and monitor their children’s posture while studying […]. Children often find it difficult to challenge the decisions of their parents […] In Chinese families, questioning the choice of a parent is often seen as disrespectful.” In addition, within Chinese culture, where hospitality is highly valued, guests may either be unaware that they are being monitored or feel uneasy knowing that smart devices are present, which could be perceived as a breach of trust and courtesy.
Additionally, P9-UX, P18-LE, and P20-LE noted the growing use of smart home devices to monitor domestic workers for security or deterrence purposes, raising ethical concerns about their privacy: “Privacy and security settings are typically designed to protect device owners, often overlooking the comfort of domestic workers who lack the power to consent to being monitored by their employers’ surveillance cameras.” Further, P12-PM observed that purchasing and installing smart home devices often mirrored existing household gender roles, stating: “Gender dynamics often shape buying decisions in typical Chinese households: When a woman suggests buying a device, she may seek input from other family members, but if a man, usually the father or husband, disagrees, the purchase often does not happen.”
@@
Law enforcement vs. companies. Some participants noted that their companies usually required legal documents, such as a search warrant, to authorize police access to user data, in accordance with PIPL (Articles 34, 35). P5-UX added that they generally avoided granting access to data unless it was essential for an investigation, sometimes obtaining the consent of the homeowner beforehand. However, some participants noted that certain companies allowed police access to all user registration and facial data collected by smart home cameras in specific emergency situations. For example, P13-UX highlighted: “Government or public security departments have special privileges to access user facial data, names, and other personal information. They have the authority to force us to provide this data and can also retrieve it through internal channels, bypassing the restrictions our teams put in place.”
@@ 4.3.1 Technical interventions
To mitigate privacy concerns and comply with laws and regulations, many participants described anonymizing or blurring image data prior to processing. As P9-UX explained, “We don’t capture real images directly—sensitive data, like nude bodies, is processed by blurring privacy zones detected by smart cameras.” Furthermore, several company guidelines imposed strict procedures for managing data, especially when it involved sensitive items such as biometric records. For example, P11-PM and P24-EN noted that camera footage was encrypted prior to transmission, and their team built specialized decoders to retrieve those recordings. P9-UX and P17-EN described employing a data-minimization approach, with P9-UX explaining: “We only send and save posture data, like simplified 2D images resembling stick figures, instead of actual scene details.” Moreover, many participants mentioned that their companies followed data deletion practices, often in accordance with regulatory mandates. For example, P4-PM, P16-PM, and P19-LE noted that Chinese regulatory bodies required companies to respect users’ rights to have their data erased. P5-UX added: “Given the large volume of data and the high cost of storage, we typically delete video footage after it has been streamed in real time, unless users have subscribed to a storage service.” P11-PM, P21-EN, and P22-EN also mentioned that users had the option to delete their cloud-stored data within 30 days, after which the system would automatically remove it.
Most participants highlighted that their companies imposed strict controls on sensitive user information, allowing only authorized personnel to access it. For example, P12-PM explained: “Our software developers have different levels of access. For example, a developer with a user account can view linked devices, whereas database and server developers are limited to operational queries, such as tracking user numbers or feature usage, without accessing individual user data. Once a feature is deployed, it is controlled by the user, and software developers no longer need to maintain or access it.”
However, some participants, especially UX designers, expressed a lack of clarity regarding how smart home devices managed and processed data, often citing limited access and strict company policies as the reason. As P6-UX stated: “We often lack a full understanding of data collection and processing because our design team operates separately from the compliance team.” P18-LE and P23-LE also underscored this issue from the compliance team’s perspective, noting limited cross-team communication.